What is JIT? a. Just in time b. Justin Timberlake c. Jobs in transit d. Just...

  1. What is JIT?

    a.

    Just in time

    b.

    Justin Timberlake

    c.

    Jobs in transit

    d.

    Just inventory transferrable

5 points

Question 15

  1. Lean accounting is

    a.

    A costing method that focuses on value streams, as distinguished from individual products or departments, thereby eliminating waste in the accounting process

    b.

    A costing method that focuses on value streams undistinguishable from individual products or departments

    c.

    A method that results in increases to waste in the accounting process

    d.

    A six sigma process focused only on improving the manufacturing processes of medium to large businesses

5 points

Question 16

  1. Chuck has been told that if he invests $500,000 on a solar system he can save $130,000 a year. The lifespan of the panels he is being sold are 10 years at which time they are rendered useless. The required rate of return is 10%.   Ignore any tax and assume all cash flows occur at year end except for the initial investment amounts. Calculate the internal rate of return using the interpolation method. Round all results in each calculation to the thousandths.

    a.

    19.892%

    b.

    22.640%

    c.

    10.271%

    d.

    10.190%
Answer #1

Solution

(14)
Answer: Option (a) Just-in-Time

(15)
Answer: Option (a) A costing method that focuses on value streams, as distinguished from individual products or departments, thereby eliminating waste in the accounting process

(16)
Initial Investment = $ 500,000
Annual Savings = $ 130,000
Lifespan = 10 Years

At Internal Rate of Return (IRR), Discounted Cash Flow = Initial Investment

Therefore, by making trial and error calculation,
Taking IRR at 22%, Discounted Cash Flow = $ 130,000 x Present Value Annuity Factor of $ 1 @ 22% for 10 Years
= $ 130,000 x 3.9232
= $ 510,016
Again, taking IRR at 23%, Discounted Cash Flow = $ 130,000 x Present Value Annuity Factor of $ 1 @ 23% for 10 Years
= $ 130,000 x 3.7993
= $ 493,909

Therefore, the internal rate of return should be in between 22% and 23%. Therefore, assuming IRR = X, value using interpolation will be,

(X - 22) / (23 - 22) = (500,000 - 510,016) / (493,909 - 510,016)
Or, X - 22 = 10,016 / 16,107
Or, X - 22 = 0.622 (Approx.)
Or, X = 22.622

Due to approximation error, the IRR have been computed as 22.622%.

Answer: Option (b) 22.640%

Present Value Annuity Table for 10 Years
10% 6.1446
11% 5.8892
12% 5.6502
13% 5.4262
14% 5.2161
15% 5.0188
16% 4.8332
17% 4.6586
18% 4.4941
19% 4.3389
20% 4.1925
21% 4.0541
22% 3.9232
23% 3.7993